OpenStudy (anonymous):

Find the slope of the line passing through the points (-7,9) and (-7,3)

OpenStudy (anonymous):

(3-9)/(-7-(-7))=-6/0=undefined.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -7}}\quad ,&{\color{blue}{ 9}})\quad &({\color{red}{ -7}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)